# Evolving from Single-modal to Multi-modal Facial Deepfake Detection: A Survey

### v1

@@ 3.2.2 Text-Visual
Liu et al. [168] incorporate frequency domain analysis into RGB image evaluations using discrete wavelet transforms to capture forgery artifacts more effectively. Their approach integrates image and frequency data through a specialized encoder and a unified decoder, simplifying the optimization process while improving forgery detection.
@@
3.3 Trustworthy Deepfake
Despite considerable advances in deepfake detection, from single-modal to multi-modal approaches, concerns persist regarding the robustness and trustworthiness of these detectors. For example, a recent study by Abdullah et al. [177] highlights significant vulnerabilities in current detection methods. The study indicates that attackers can easily tailor generative models to circumvent traditional defenses. Moreover, the misuse of widely deployed vision foundation models allows for the creation of adversarial deepfakes through subtle manipulations of image semantics, thereby undermining the effectiveness of established detection techniques. In the following subsections, we will explore recent developments in deepfake detection, focusing on adversarial attacks, backdoor attacks, and defensive strategies.
Adversarial Attack Adversarial attacks in deepfake detection aim to deceive manipulation detectors by subtly altering original images across various domains—pixel space, frequency domain, and latent space. Carlini et al. [178] highlight how adversarial strategies can significantly lower the accuracy of deepfake detectors. Supporting studies [179, 180, 181] confirm these vulnerabilities, emphasizing the need for robust detection mechanisms.
In the pixel domain, Neekhara et al. [182] explore how adversarial attacks target specific facial features, while Liao et al. [183] develop the Multi-Layers Key Semantic Region Selection method to target essential pixels. Zhu et al. [184] and Liu et al. [185] introduce less noticeable perturbations using frequency and spatial features to streamline the application across various images. In [186], Huang et al. introduce DeepNotch, a method to create realistic and undetectable DeepFakes by reducing artifact patterns without compromising image quality, using implicit spatial-domain notch filtering. For frequency-based attacks, Jia et al. [187] and Chen et al. [188] focus on disrupting frequency domain properties to fool detection systems, utilizing cross-task perturbed initialization to enhance attack efficacy. In [189], Wu et al. introduce TraceEvader, a novel evasion attack that manipulates high-frequency traces and introduces adversarial blur into low-frequency components. In latent space manipulations, Meng et al. [190], Shamshad et al. [191], and Li et al. [192] leverage the latent spaces of generative models to enhance the stealth and effectiveness of their attacks, merging semantic manipulations with adversarial perturbations to bypass forensic classifiers.
In a departure from traditional 2D adversarial attacks, Shahriyar’s studies [193] and the innovative ”AdvHeat” by Wang et al. [194] explore new dimensions such as video content dynamics and 3D adversarial face views.
Backdoor Attack Unlike adversarial attacks that manipulate input at runtime, backdoor attacks embed corrupted samples in training datasets. These samples appear legitimate, but contain Trojans that, when triggered, manipulate the model’s output to evade detection. Triggers can be placed in the pixel space [195, 196] or latent space [197], and can be selected from being manually specified [195] to stealthy [197, 196].
Sun et al. [195] propose to craft effective triggers by exploiting the vulnerabilities of deepfake detectors and select impactful samples with a Filtering-and-Updating Strategy. Han et al. [197] use an ”analysis-by-synthesis” approach to embed natural triggers in the latent space, involving Model Discrimination and Data Distribution to create independent poisoned samples.
Liang et al. [196] focus on embedding inconspicuous triggers, limiting perturbations to critical facial areas and employing a scalable trigger generator that creates transformation-sensitive patterns. This method enhances the stealth of triggers and addresses their sensitivity to image transformations, significantly improving the subtlety of backdoor attacks.
Discrepancy Minimization This approach focuses on reducing the detectable differences between genuine and manipulated images across various data representation domains, including pixel space, frequency space, and feature space.
In pixel space, Huang et al. [198] develop FakePolisher, which reduces artifacts by projecting images into a low-dimensional subspace. Their subsequent FakeRetouch [199] enhances image realism through strategic noise addition and image filtering. Similarly, Li et al. [200] and He et al. [201] improve facial data reconstruction by searching adversarial latent codes.
In the frequency domain, Dong et al. [202] balance spectral power distributions to make discrepancies less noticeable, while Lee et al. [203] introduce Spectrum Translation for refining image generation spectra.
For feature space discrepancies, Hou et al. [204] minimize differences by applying natural degradations in an adversarial manner. Ivanovska et al. [205] employ Denoising Diffusion Models to challenge detection systems, focusing on reconstructing images with specific denoising steps for robust testing. Additionally, Liu et al.’s TR-Net [206] efficiently removes intrinsic deepfake traces through coordinated generators and multiple discriminators, enhancing the method’s effectiveness across various detection systems.
Defense Strategies To enhance the resilience of deepfake detection systems against increasingly sophisticated attacks, researchers have developed robust defense strategies. Khan et al. introduce Adversarial Feature Similarity Learning (AFSL) [207], which improves detection robustness by optimizing feature similarity and implementing strategic regularizations. Additionally, Pinhasov et al. utilize Explainable Artificial Intelligence (XAI) to create interpretability maps, aiding in the robust classification of adversarial attacks [208]. The approach proposed in [209] improves the robustness of deepfake defenses through a destruction-aware constraint during training. [210] introduces FaceGuard that enhances defense against unseen adversarial attacks by generating various perturbations that evade detection, while concurrently learns to identify and remove these perturbations from images. Ensemble methods have also been prominent, with multiple researchers employing these strategies to capitalize on the diverse strengths of combined models, significantly enhancing detection accuracy and robustness against adversarial perturbations [211, 212, 213, 214]. These methods underscore a critical shift towards integrating advanced computational techniques to safeguard against the evolving landscape of digital forgery.
@@ 4 Challenges and Future Directions
As discussed in the last section, deepfake detection has achieved significant progress in recent years. However, various challenges persist due to the complexity of deepfake detection. In this section, we address these challenges and present potential avenues for future research.
Performance Generalization Across Scenarios Current deepfake detection methods show limited effectiveness in generalized deepfake detection, including across different datasets, manipulation methods, and post-processing techniques. The variations across these factors can result in significant statistical distribution differences between the training and testing sets. As pointed out in previous research on domain adaptation [215, 216] and domain generalization [217], ignoring these differences may lead to a dramatic performance drop when applying deepfake detectors in practical scenarios. Additionally, adversarial attacks pose another challenge on the robustness of models. Considering these challenges and the need for practical applicability, developing specific methods for generalized and robust deepfake detection becomes a promising and important direction.
@@
Real-Time Analysis Deepfake detection, especially multi-modal detection, involves manipulation across different modalities and often occurs in the format of video/audio streams. In such cases, analysis and prediction can only be made after processing all streams from various modalities, which can be time-consuming. However, the potential risk introduced by the spread of deepfakes requires a real-time analysis and response [218]. Current methods focus more on prediction accuracy than efficiency, highlighting the need for further innovation and additional efforts to achieve real-time detection.
Expanding Deepfake Detection with LLMs A recent study [17] has explored using large language models (LLMs) for generation and detection. We believe that integrating LLMs into deepfake detection can be a promising direction, leveraging their powerful capabilities in detection and explainability. For example, LLMs can analyze subtleties in the text modality that may accompany deepfake videos, such as inconsistencies between spoken words and lip movements. Exploiting these discrepancies can enhance the analysis. Furthermore, LLMs can automatically generate explanations about detected deepfakes, detailing the reasons for the classification and providing evidence, which is crucial for trust and verification purposes.
@@ 5 Conclusion
This paper provides a comprehensive and up-to-date overview of deepfake detection, summarizing methods accepted by top-tier conferences and journals over the past three years. We first formally define the deepfake detection problem within our scope, followed by detailed datasets and evaluation metrics for deepfake detection, covering various modalities and generation models. Next, we systematically explore the mainstream deepfake detection methods and categorize them based on their characteristics. We believe that our careful categorization will help researchers quickly grasp the development trends in this field. Finally, we delve into the challenges and potential future directions in facial deepfake detection.

### v3

@@ 3.2.2 Text-Visual Deepfake Detection
Liu et al. [188] incorporate frequency domain analysis into RGB image evaluations using discrete wavelet transforms [201] to capture forgery artifacts more effectively. Their approach integrates image and frequency data through a specialized encoder and a unified decoder, streamlining the optimization process. Additionally, the authors utilize the unified decoder featuring symmetric cross-modal interaction modules for gathering modality-specific forgery information and a fusing interaction module to aggregate both visual and textual features.
@@ 4 Challenges and Future Directions
As discussed in previous sections, deepfake detection has achieved significant progress in recent years. However, various challenges persist due to the complexity of deepfake detection. In this section, we discuss these challenges and present potential avenues for future research.
Performance Generalization Across Scenarios Current deepfake detection methods show limited effectiveness in generalized detection across different datasets, manipulation methods, and post-processing techniques. Variations in these factors can cause significant statistical distribution differences between training and testing sets. Previous research on domain adaptation [202] and domain generalization [203] indicates that ignoring these differences can lead to dramatic performance drops in practical scenarios. Additionally, adversarial attacks challenge the robustness of the models. Considering these challenges, developing specific methods for generalized and robust deepfake detection is a crucial direction for future research.
@@
Expanding Deepfake Detection with LLMs A recent study [17] has explored using LLMs for generation and detection. We believe that integrating LLMs into deepfake detection can be a promising direction, leveraging their powerful capabilities in detection and explainability. For example, LLMs can analyze subtleties in the text modality that may accompany deepfake videos, such as inconsistencies between spoken words and lip movements. Exploiting these discrepancies can enhance the analysis. Furthermore, LLMs can automatically generate explanations about detected deepfakes, detailing the reasons for the classification and providing evidence, which is crucial for trust and verification purposes.
Privacy-Aware Facial Deepfake Detection Ensuring data privacy in facial deepfake detection is critical due to the sensitive nature of the data and increasing privacy regulations. Most prior works use centralized training, aggregating facial data and increasing the risk of data leakage. Future research should explore privacy-aware methods like federated learning [204]. Federated learning allows training across decentralized devices without data exchange, preserving user privacy [205, 206, 207]. These approaches can enhance compliance with privacy regulations while maintaining high detection performance.
@@ 5 Conclusion
This paper provides a thorough review of recent developments in deepfake detection, highlighting research from leading conferences and journals over the past three years. It starts by clearly defining the scope of deepfake detection, followed by an in-depth discussion of datasets and evaluation metrics that address various modalities and generative models. The study methodically reviews and categorizes the principal deepfake detection methods, aiming to help researchers quickly understand current trends. The paper concludes with a discussion on the challenges and prospective future directions in facial deepfake detection.
Acknowledgments Dr. Joey Tianyi Zhou is supported by the National Research Foundation, Prime Minister’s Office, Singapore, and the Ministry of Digital Development and Information, under its Online Trust and Safety (OTS) Research Programme (MCI-OTS-001) and SERC Central Research Fund (Use-inspired Basic Research). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore, or the Ministry of Digital Development and Information.